LaylaSanaii:

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 11 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section? 1- 6 over 20 2- 3 over 20 3- 9 over 20 4- 17 over 20

Tonycoolkid21:

am pro

LaylaSanaii:

@tonycoolkid21 wrote:
am pro
huh?

Tonycoolkid21:

why you helping me with my work

LaylaSanaii:

@tonycoolkid21 wrote:
why you helping me with my work
trying to be nice

Tonycoolkid21:

k

xhikari:

okay so there are 20 sections total, they give us the amount of red and blue but not green.. 6+11=17 so what is leftover from 20-17 = ?

LaylaSanaii:

3

LaylaSanaii:

3 over 20 ?

xhikari:

therefore the probablity is 3/20.. when we add 6/20 + 11/20 + 3/20 = 20/20

xhikari:

yes great work! :)

LaylaSanaii:

@xhikari wrote:
yes great work! :)
thx

xhikari:

@laylasanaii wrote:
@xhikari wrote:
yes great work! :)
thx
np <3